OpenStudy (anonymous):

a. What is the speed at t=20.0sec? b. Find the distance traveled in the first 20 seconds

OpenStudy (anonymous):

here's the graph

OpenStudy (amistre64):

what is teh acceleration at t=20?

OpenStudy (amistre64):

lol, read it a little off ... start at the begining; let v = at when a=0, then v remains constant

OpenStudy (amistre64):

something akin to that ... but my last leg might be off 

OpenStudy (amistre64):

v = 2t, from 0 to 10 v = 2(10), from 10 to 15 v = -3t+c, from 15 to 20 v = -3(20)+c from 20 onwards

OpenStudy (anonymous):

What does C represent? And for the second part of my question, I got the answer 237.5, and the website I'm using to submit my assignment says it is within 10% of the correct answer, so I'm having trouble fixing my mistake. I found the area underneath each part of hte graph, 100 from 0-5 sec, 100 from 10 to 15 sec, and 37.5 from 15-20sec

OpenStudy (amistre64):

some concepts first: we agree that v = at ?

OpenStudy (amistre64):

c just represents the shift by starting at 15 instead of 0

OpenStudy (amistre64):

we could just as well consider this in three parts; v = 2t, from 0 to 10 v = 2(10), from 0 to 5 v = 20 - 3t, from 0 to 5

OpenStudy (amistre64):

each part is calibrated to start the count anew is all

OpenStudy (amistre64):

the displacement is just the areas of the 3 different parts of velocity triangle: 10x20 rectangle: 20x5 rectangle-ish: (20+5)/2 x 5

OpenStudy (amistre64):



OpenStudy (amistre64):

182.5 displacement seems good to me

OpenStudy (amistre64):

forgot how to do math .... lol

OpenStudy (amistre64):

100+100+5(25)/2 = 262.5

OpenStudy (anonymous):

ok that's what I got lol!

OpenStudy (anonymous):

Can we go back to the speed at t=20?

OpenStudy (amistre64):

t=20 is the end of the third leg of this which i simplified as: v = 20 - 3t, from 0 to 5 since it only lasts 5 seconds and it start at 20

OpenStudy (anonymous):

So i it would be 5? I was thinking you would do v+at with a=-3 and t=5

OpenStudy (amistre64):

the trip starts at v=0 a=2, from 0 to 10 seconds; v = 0 to 20 for the next 5 seconds they are just cruising along at 20 .. now they start to de-accelerate by -3: 20 - 3t for the next 5 seconds; ending in: 20-3(5) = 5 for a final velocity

OpenStudy (amistre64):

poor choice of words ... de-accelerate at 3 or accelerate at -3 ...

OpenStudy (anonymous):

Oh! I understand now thank you!

OpenStudy (amistre64):

youre welcome